# A Comprehensive Survey of Foundation Models in Medicine

### v1

@@ 3 Overview of Foundation Models
3.1 Learning architecture
The underlying architecture of FMs is typically based on transformers capable of capturing long-range dependencies and contextual relationships within data (Section 2.2). The main architectures are encoder-decoder, encoder only, and decoder only are explained below:
3.1.1 Encoder-decoder
@@ 3.1.3 Decoder-only
Decoder-only models specialize in the creative generation of coherent sequences. They predict the next token when given a sequence of tokens, which results in the generation of contextually meaningful outputs. Many state-of-the-art models, such as BLOOM [81], PaLM [82], and open pre-trained transformers [83], are based on decoder-only architecture. GPT [15] models, such as GPT-2 [60], GPT-3 [29], and the more recent GPT-4 [84], are decoder-only models pre-trained on large-scale unsupervised text data.
3.2 Learning algorithms
@@
SSL enables FMs to learn rich contextual representations from a large amount of unlabeled data. The learned representations can then be fine-tuned for specific downstream tasks. SSL tackles the challenges associated with expensive, biased annotation and narrowly applicable supervised learning by providing a more robust and efficient approach for generalizing across downstream tasks [85]. SSL methods that require supervised fine-tuning steps with manually labeled data have limited usability, especially for medical datasets, which are difficult to collect and annotate [86]. Therefore, zero-shot learning-based methods are used instead, which do not require explicit manual or annotated labels during training [86]. Zero-shot learning-based methods include the GPT models [15, 60, 29, 84], CLIP [57], PaLM[82], and SLIP [68].
3.3 Human feedback reinforcement learning
@@
FM models that learn from human feedback tend to demonstrate improved performance and can be generalized to other domains, unlike supervised learning [87]. For instance, Stiennon et al. [87] used the GPT-3 model with 6.7 billion parameters to predict human-preferred summaries from large-scale data and used the model as a reward function to fine-tune a summarization policy through RL. Specifically, for Reddit posts [88], multiple summaries were sampled from various sources including the current policy, initial policy, original reference summaries, and baselines. The summaries were then evaluated by human labelers and a reward model was trained to predict the likelihood that one summary was preferred over the other, based on human judgments.
3.4 Prompting Foundation Models
@@
Prompting FMs are directed to perform a specific task to enhance their effectiveness and relevance [35]. In promoting FMs, users provide a prompt or a series of prompts that guide the model’s response generation process. It aids the model to generate a response that is more aligned with the user’s expectations. Incorporating additional context into the training and evaluation prompting FMs aids in improving the model’s performance by providing relevant information and guiding its decision-making process. For instance, GPT-based application [29, 84] possesses strong language understanding capabilities that can be prompted with specific textual inputs, such as questions. These can be further fine-tuned for tasks like text generation, sentiment analysis, or document summarization. Furthermore, prompts can be used to specify the desired task such as image classification or caption generation that allows the model to focus its attention and resources accordingly. Models such as CLIP [57] which is trained on both image and text encoders using a contrastive pre-training objective rely on textual prompts for supervision. Furthermore, CLIP variants such as MaskCLIP [69] and RegionCLIP [89] are based on object detection and segmentation to align image regions with textual descriptions. We refer readers to [35] for further details about prompting models.
4 Flagship Foundation Models
This section highlights the key foundation models and their variants.
4.1 BERT
@@
BERT [16] is a transformer-based model that was developed for language representation through training on a large corpus of textual data. It is a pre-trained bidirectional transformer model that utilizes contextual information within sequences from unlabeled data through the use of masked language modeling (MLM). The BERT frameworks consist of two main stages: pre-training and fine-tuning. The pretraining stage consists of two tasks: MLM and next sentence prediction (NSP). For MLM, 15% of input tokens are randomly masked, and the model predicts these masked tokens. However, [MASK] tokens are not present during fine-tuning, causing a mismatch between pre-training and fine-tuning phases. To address this, the authors implemented a strategy during pre-training where [MASK] tokens were utilized 80% of the time, and a random token 10% of the time, and the unchanged token was used 10% of the time during pre-training. The authors showed that existing techniques were unable to determine the relationship between multiple sentences and were only focused on individual sentences. Therefore, NSP was used to predict the next sentence that achieved over 97% accuracy for the binary classification task (next sentence vs not next sentence).
@@
During fine-tuning, a self-attention mechanism is used for downstream tasks with a limited amount of labeled data. Originally, there were two BERT variants (BERTBASE and BERTLARGE) distinguished by several transformer blocks, hidden layers, and self-attention heads. BERTBASE had 110M parameters while BERTLARGE consisted of 340M parameters. However, following the success of BERT, multiple variants of BERT models were introduced. For instance, a more efficient and robust model called RoBERTa (robustly optimized BERT pretraining approach) [17] surpassed the BERT model using a few modifications to the original BERT. For example, the RoBERTa is trained for longer (500K steps) on a large dataset of over 160 GB of English-language corpora. Furthermore, the [MASK] token was applied dynamically which creates a new masking pattern each time at the input to the model. In addition, the NSP loss module was removed.
@@
Although BERT has been reasonably successful, its computation complexity and communication overhead are major limitations to adopting such models at scale [90]. Therefore, models inspired by BERT were developed to achieve BERT-level performance while reducing the training time and model size. For instance, Sanh [91] proposed the lightweight DistilBERT, which is a smaller and faster language representation model. DistilBERT is based on the distillation knowledge compression technique [92], which is a smaller model (student) and trained to reproduce a larger model (teacher). The number of layers in DistilBERT is reduced by half and the architecture has been optimized [91]. Furthermore, the triple loss was used during pretraining: masked language modeling loss, cosine embedding loss, and distillation loss. The training objective in distilBERT included the combination of masked language modeling loss, distillation loss, and cosine embedding loss to align student and teacher hidden state vectors. The proposed model was nearly 40% smaller in size and 60% faster than BERT with comparable performance on various downstream tasks.
@@
A Lite BERT (ALBERT) [90] introduced factorized embedding parameterization and cross-layer parameter sharing that incorporated BERT. Furthermore, the NSP module was replaced with a self-supervised loss for sentence order prediction to improve its performance. The proposed model was 1.7 times faster, used 18 times fewer parameters, and achieved better performance than BERT. The use of BERT-based models have been deployed in healthcare for various applications with great success. A model based on a denoising autoencoder known as BART [18] has been developed for sequence-to-sequence modeling and shows promising performance. However, it is computationally expensive, with 10% more parameters than BERT.
4.2 GPT
@@
The ChatGPT we interact with today originated from extensive research trials based on a GPT model. Radford et al. [15] proposed a transformer-based GPT model that utilizes an unsupervised approach during pre-training and supervised fine-tuning for natural language understanding tasks. GPT was trained on the BooksCorpus dataset [93] for unsupervised pre-training. Then, the parameters learned at the pre-training stage were used for a target task with a manually annotated dataset in a supervised manner. A separate work by Radford et al. [60] presented a GPT-2 language model with a zero-shot setting for multiple downstream tasks without the need for any major modifications. GPT-2 was trained on a large amount of data — WebText — that contains over 8 million cleaned documents with more than 40 GB of text. GPT2 performed well on multiple tasks, including reading comprehension, summarization, translation, and question-answering. However, although it performed better than baselines, GPT-2 still could not be applied practically as the performance was still not promising. Therefore, a more robust model, GPT-3, with 175B parameters, was presented and performed well on various downstream tasks. GPT-3 uses the same architectures as GPT-2 except for the implementation of alternating dense and locally banded sparse attention patterns in the transformer layers. It was trained on a large amount of data with 570 GB (300 billion tokens) of plain compressed filtered text. The performance of GPT-3 for various downstream tasks surpasses that of existing techniques when tested using zero-shot, one-shot, and few-shot learning scenarios. One key solution to reduce untruthful or irrelevant outputs is to incorporate human feedback, such as in InstructGPT [94]. InstructGPT uses reinforcement learning from human feedback (RLHF) to fine-tune the GPT-3 model on human data and reduce the number of parameters to only 1.3B. Another more sophisticated language model known as ChatGPT  [^3] https://chat.openai.com/ or GPT3.5 provides more user-friendly, interactive, and context-aware responses and conversational experience. It follows a similar refined architecture to that of InstructGPT [94]. There are a range of healthcare applications that utilize ChatGPT and these will be covered in later sections. ChatGPT is limited to text inputs, but this limitation was addressed in the GPT-4 [84] model, which accepts images in addition to text inputs while producing text outputs. However, the details of GPT4 are not publicly available and a major issue with such language models is hallucinations [84].
4.3 DALL-E
@@
DALL-E [95] is a transformer-based text-to-image generative model that employs a two-stage training procedure. In the first stage, each RGB image is compressed into a 32$\times$32 grid of image tokens using a discrete variational autoencoder (dVAE). It bypasses the direct use of pixels, significantly reducing memory consumption. Furthermore, short-range dependencies between pixels are prioritized during training [96]. Therefore, in the second stage, an autoregressive transformer models the joint distribution over text and image tokens. The training maximizes the evidence lower bound (ELB) on the model’s joint likelihood over images, captions, and tokens. The proposed model was trained on a dataset of 250 million text-image pairs similar to JFT-300M [97] collected from the internet.
An extension of DALL-E known as DALL-E-2 [98] has also been presented, which can generate high-resolution images based on text prompts and existing images. It uses CLIP-based robust representation to enhance its image generation. DALL-E-2 first generates an image embedding induced by the text input using CLIP. Then, this generated image embedding is used as an input of the diffusion model to create an image. The model is trained using an image ($x$)-caption ($y$) pair. The proposed generative model comprises a prior ($P(z_{i}|y)$) producing CLIP image embeddings $z_{i}$ conditioned on $y$. Meanwhile, the decoder ($P(x|z_{i},y)$) generates images conditioned on CLIP image embeddings. These components are stacked to yield the generative model ($P(x|y)$).
4.4 Llama
@@
LLaMa [99] is trained on publicly available datasets with over 1.4 trillion tokens and is based on existing large pre-trained models with key modifications. For instance, pre-normalization from GPT 3 was adopted to improve its training stability, but unlike in GPT-3, normalization was performed at the input of each transformer sub-layer. The SwiGLU [100] activation function from the PaLM model [82] replaced the ReLU non-linearity activation function. The absolute positional embeddings were replaced with rotary positional embeddings (RoPE) [101] from the PaLM model [82] at each layer. The training optimization included the use of the AdamW optimizer [102] and efficient implementation techniques for causal multi-head attention, backward passing, and memory usage reduction. LLaMA has 65B parameters several times more than GPT-3, and performed well in various downstream tasks. A more refined version called LLaMA 2 [103], with 70B parameters in addition to Llama 2-Chat, is designed for dialogue applications and performance surpasses open-source chat models on various benchmarks [103]. LLaMA2 incorporates increased context length and the implementation of grouped-query attention (GQA) to enhance inference scalability for larger models.
4.5 Stable Diffussion
@@
Diffusion models (DMs) [104] can synthesize high-quality images but demand significant computational resources during training. Furthermore, they require costly evaluations in pixel space. Therefore, an efficient model called the stable diffusion model [105] explicitly separates the compressive and generative learning phases. The proposed model employs an autoencoder that generates a lower-dimensional representation space and significantly reduces computational complexity. This enhances the computational efficiency of sampling, leveraging the inductive bias of UNet-based diffusion models for spatially structured data and providing a multipurpose compression model applicable to various generative tasks. The proposed diffusion model demonstrates competitive performance in tasks such as unconditional image synthesis, inpainting, and super-resolution, significantly reducing computational costs compared to pixel-based DMs. An improved version of the stable diffusion model, known as SDXL [106], incorporates a larger UNet backbone (2.6B parameters in the UNet), additional attention blocks, and a broader cross-attention context. Furthermore, it uses a refinement model, enhancing visual fidelity through post-hoc image-to-image techniques.
4.6 Industrial pretrained models
@@
Industrial-scale pre-trained models are extensively employed across a spectrum of industrial applications. Here we briefly mention some state-of-the-art models widely adopted in various industries including healthcare. LaMDA (Language Models for Dialog Applications) [107] is a transformer-based model with 137B parameters tailored for dialog applications. It addresses two challenges faced by language models: safety and factual grounding. The safety enhancements involve filtering responses using a classifier fine-tuned with crowd worker-annotated data, aligning model behavior with human values based on Google AI principles [^4] https://ai.google/responsibility/principles/. Factual grounding is achieved by enabling the model to consult external knowledge sources to ensure responses are grounded in verifiable information.
@@
A more robust model than GPT-3 known as GLaM (Generalist Language Model) [108] has been proposed by Google, which has 1.2 trillion parameters. It utilizes sparsely activated mixture-of-experts (MoE) architecture [109] in which every other transformer layer replaces the feed-forward component with an MoE layer. GLaM achieved better performance on 29 NLP tasks in nearly all zero-shot, one-shot, and few-shot learning scenarios, consuming only one-third of the energy of GPT-3. A more refined version, the Pathways Language Model (PaLM) [82], with 540 billion parameters, has also been presented by Google. It is a decoder-only-style transformer model with multiple modifications, such as SwiGLU activations [100] for MLP intermediate layers, parallel formulation for a faster training speed, multi-query attention, and RoPE embeddings [101] for improved performance on long sequences. PaLM has achieved state-of-the-art performance in reasoning tasks and shows proficiency in multilingual tasks and source code generation. A more efficient model, PaLM 2 [110], was recently developed by Google and performs better than PaLM on various downstream tasks, with enhanced multilingual, reasoning, and coding capabilities. It is utilized at various sizes in 25 Google products, including Bard  [^5] https://bard.google.com/, the health-focused Med-PaLM [111], Med-Plam 2 [112], and the security-focused Sec-PaLM [^6] https://cloud.google.com/security/ai/. SenseNova [^7] https://www.sensetime.com/en/news-detail/51166397?categoryId=1072, developed by SenseTime, is a large-scale FM that offers various API interfaces and services specifically designed for enterprise customers. It is widely adopted in various industries, mainly in China, such as in the automotive sector, where SenseNova played a key role in the creation of the BEV (Bird’s-Eye-View) system that can recognize 3,000 types of objects. Furthermore, SenseNova introduced the “SenseChat” [^8] https://chat.sensetime.com/wb/login model, which is similar to ChatGPT but specifically designed for use in China.
Table 2: Flagship Foundation Models
@@
[16] 2018 BERTBASE 110M [93] [^9] BooksCorpus, English Wikipedia NLP
[16] 2018 BERTLARGE 340M [93] , English Wikipedia NLP
[17] 2019 RoBERTa 355M [93], English Wikipedia, CC-News, Opne Web Text [113], stories [114] NLP
[91] 2019 DistilBERT 66M [93], English Wikipedia NLP
[90] 2020 ALBERTBASE 12M [93], English Wikipedia NLP
[90] 2020 BERTLARGE 18M [93] , English Wikipedia NLP
[90] 2020 BERTxLARGE 60M [93], English Wikipedia NLP
[90] 2020 BERTxxLARGE 235M [93], English Wikipedia NLP
@@
[60] 2019 GPT-2 1.5B WebText NLP
[29] 2020 GPT-3 175B Common Crawl (filtered), WebText2, Books1, Books2, Wikipedia NLP
@@
[94] 2022 InstructGPT 1.3B Common Crawl (filtered), WebText2, Books1, Books2, Wikipedia NLP
@@
[99] 2023 LLaMA 65B CommonCrawl, C4 [19], Github, Wikipedia, Books, ArXiv, StackExchange NLP
@@
[106] 2023 SDXL 3.5B Not reported Image synthesis
@@
[108] 2022 GLaM 137B Filtered Webpages, Wikipedia, Conversations, Forums, Books, News NLP tasks

### v2

@@ Appendix B Overview of Foundation Models
B-A Learning architecture
The underlying architecture of FMs is typically based on transformers that can capture long-range dependencies and contextual relationships within data (Section A-B). The main architectures are encoder-decoder, encoder only, and decoder only, are explained below:
B-A1 Encoder-decoder
@@ B-A3 Decoder-only
Decoder-only models specialize in the creative generation of coherent sequences. They predict the next token when given a sequence of tokens, which results in the generation of contextually meaningful outputs. Many SOTA models, such as BLOOM [300], PaLM [301], and open pre-trained transformers [302], are based on decoder-only architecture. GPT [12] models, such as GPT-2 [57], GPT-3 [28], and the more recent GPT-4 [303], are decoder-only models pre-trained on large-scale unsupervised text data.
B-B Learning algorithms
@@
SSL enables FMs to learn rich contextual representations from unlabeled data that can be fine-tuned for specific downstream tasks. SSL tackles the challenges associated with expensive, biased annotation and narrowly applicable supervised learning by providing a more robust and efficient approach for generalizing across downstream tasks. SSL methods that require supervised fine-tuning steps with manually labeled data have limited usability, especially for medical datasets, which are difficult to collect and annotate [129]. Therefore, zero-shot learning-based methods are used instead, which do not require explicit manual or annotated labels during training [129]. Zero-shot learning-based methods include the GPT models [12, 57, 28, 303], CLIP [56], PaLM[301], and SLIP [304].
B-C Human feedback reinforcement learning
@@
FM models that learn from human feedback tend to demonstrate improved performance and can be generalized to other domains, unlike supervised learning [305]. For instance, Stiennon et al. [305] used the GPT-3 model with 6.7 billion parameters to predict human-preferred summaries from large-scale data and used the model acts as a reward function to refine a summarization policy through RL. Specifically, for Reddit posts [306], multiple summaries were sampled from various sources. The summaries were then evaluated by human labelers and a reward model was trained to predict the likelihood that one summary was preferred over the other, based on human judgments.
B-D Prompting Foundation Models
@@
Prompting FMs are directed to perform a specific task to enhance their effectiveness and relevance [34]. In promoting FMs, users provide a prompt or a series of prompts that guide the model’s response generation process. It aids the model to generate a response that is more aligned with the user’s expectations. Incorporating additional context into the training and evaluation prompting FMs aids in improving the model’s performance by providing relevant information and guiding its decision-making process. For instance, GPT-based application [28, 303] possesses strong language understanding capabilities that can be prompted with specific textual inputs, such as questions. These can be further fine-tuned for tasks like sentiment analysis, text generation, or document summarization. Furthermore, prompts can be used to specify the desired task such as image classification or caption generation that allows the model to focus its attention and resources accordingly. Models such as CLIP [56] which is trained on both image and text encoders using a contrastive pre-training objective rely on textual prompts for supervision. Furthermore, CLIP variants such as MaskCLIP [307] based on object detection and segmentation to align image regions with textual descriptions.
Appendix C Flagship Foundation Models
This section highlights the key FMs and their variants.
C-A BERT
@@
BERT [13] was developed for language representation through training on a large corpus of textual data. It is a pre-trained bidirectional transformer model that utilizes contextual information within sequences from unlabeled data through the use of masked language modeling (MLM). The BERT frameworks consist of two main stages: pre-training and fine-tuning. The pretraining stage consists of two tasks: (i). MLM, in which the model randomly masks 15% of input tokens and predicts these masked tokens. However, [MASK] tokens are not present during fine-tuning, causing a mismatch between pre-training and fine-tuning phases. To address this, the authors implemented a strategy during pre-training where [MASK] tokens were utilized 80% of the time, a random token 10%, and the unchanged token was used 10% of the time during pre-training. The authors showed that existing techniques were unable to determine the relationship between multiple sentences and were only focused on individual sentences. Therefore, NSP was used to predict the next sentence that achieved over 97% accuracy for the binary classification task (next sentence vs not next sentence).
@@
During fine-tuning, a self-attention mechanism is used for downstream tasks with limited labeled data. Multiple variants of BERT models were then introduced such as RoBERTa (robustly optimized BERT pretraining approach) [14] that surpassed the BERT mode. It is trained for longer (500K steps) on a large dataset of over 160 GB of English-language corpora. Furthermore, the [MASK] token was applied dynamically, creating a new masking pattern each time at the input to the model. In addition, the NSP loss module was removed.
@@
Although BERT has been reasonably successful, its computation complexity and communication overhead are major limitations to adopting such models at scale [308]. Therefore, models inspired by BERT were developed to achieve BERT-level performance while reducing the training time and model size. For instance, Sanh [309] proposed the lightweight DistilBERT. It is based on the distillation knowledge compression technique [310], which is a smaller model (student) and trained to reproduce a larger model (teacher). The number of layers in DistilBERT is reduced by half and the architecture has been optimized [309]. The proposed model was nearly 40% smaller in size and 60% faster than BERT with comparable performance on various downstream tasks.
@@
A Lite BERT (ALBERT) [308] was 1.7 times faster, used 18 times fewer parameters, and achieved better performance than BERT. A model based on a denoising autoencoder known as BART [15] has been developed for sequence-to-sequence modeling and shows promising performance. However, it is computationally expensive, with 10% more parameters than BERT.
C-B GPT
@@
The ChatGPT we interact with today originated from extensive research trials based on a GPT model. Radford et al. [12] proposed a transformer-based GPT model that utilizes an unsupervised approach during pre-training and supervised fine-tuning for natural language understanding tasks. GPT was trained on the BooksCorpus dataset [311] for unsupervised pre-training. Then, the parameters learned at the pre-training stage were used for a target task with a manually annotated dataset in a supervised manner. A separate work by Radford et al. [57] presented a GPT-2 language model with a zero-shot setting for multiple downstream tasks without the need for any major modifications. GPT-2 was trained on a large amount of data — WebText — contains over 8 million cleaned documents with more than 40 GB of text. GPT2 performed well on multiple tasks, including reading comprehension, summarization, translation, and question-answering. However, although it performed better than baselines, GPT-2 still could not be applied practically as the performance was still not promising. Therefore, a more robust model, GPT-3, with 175B parameters, was presented and performed well on various downstream tasks. It was trained on 300 billion tokens (570 GB) of plain compressed filtered text. The performance of GPT-3 for various downstream tasks surpasses that of existing techniques when tested using zero-shot, one-shot, and few-shot learning scenarios. One key solution to reduce untruthful or irrelevant outputs is to incorporate human feedback, such as in InstructGPT [312]. InstructGPT uses reinforcement learning from human feedback (RLHF) to fine-tune the GPT-3 model on human data and reduce the number of parameters to only 1.3B. Another more sophisticated language model known as ChatGPT  [^6] https://chat.openai.com/ or GPT3.5 provides more user-friendly, interactive, and context-aware responses and conversational experience. It follows a similar refined architecture to that of InstructGPT [312].ChatGPT is limited to text inputs, but this limitation was addressed in the GPT-4 [303] model, which accepts images in addition to text inputs while producing text outputs. However, the details of GPT4 are not publicly available and a major issue with such language models is hallucinations [303].
C-C DALL-E
@@
DALL-E [67] is a transformer-based text-to-image generative model with a two-stage training procedure. In the first stage, each RGB image is compressed into a 32$\times$32 grid of image tokens, bypassing the direct use of pixels, significantly reducing memory consumption. Furthermore, short-range dependencies between pixels are prioritized during training. Therefore, in the second stage, an autoregressive transformer models the joint distribution over text and image tokens. The proposed model was trained on a dataset of 250 million text-image pairs gathered from online sources. An extension DALL-E-2 [313] has also been presented, which can generate high-resolution images based on text prompts and existing images. It uses CLIP-based robust representation to enhance its image generation. DALL-E-2 first generates an image embedding induced by the text input using CLIP. Then, this generated image embedding is used as an input of the diffusion model to create an image. The model is trained using an image ($x$)-caption ($y$) pair.
C-D Llama
@@
LLaMa [314] is trained on 1.4 trillion tokens and is based on existing large pre-trained models with key modifications. For instance, pre-normalization from GPT 3 was adopted to improve its training stability, but unlike in GPT-3, normalization was performed at the input of each transformer sub-layer. The SwiGLU [315] activation function from the PaLM model [301] replaced the ReLU non-linearity activation function. The absolute positional embeddings were replaced with rotary positional embeddings (RoPE) [316] from the PaLM model [301] at each layer. LLaMA has 65B parameters several times more than GPT-3, and performed well in various downstream tasks. A more refined version called LLaMA 2 [317], with 70B parameters in addition to Llama 2-Chat, is designed for dialogue applications and performance surpasses open-source chat models on various benchmarks [317].
C-E LLaVA
Large Language and Vision Assistant) [318] is a large multimodal model designed to integrate vision and language for general-purpose tasks such as image understanding, captioning, and reasoning. LLaVa was built by instruction-tuning a language model using machine-generated multimodal (image-text) instruction-following data created by GPT-4. LLaVA connects a CLIP visual encoder with the Vicuna language model, fine-tuned on approximately 158,000 image-text pairs to develop its instruction-following abilities. The training process is split into two stages: first, pre-training aligns image features with the language model’s tokens, and second, end-to-end fine-tuning integrates image-text pairs to improve the model’s instruction-following capability. The model exhibits strong multimodal reasoning capabilities, achieving high performance on tasks like ScienceQA and performing well even in out-of-domain scenarios. Although LLaVA demonstrates promising performance, it struggles to interpret medical images due to complex visual patterns accurately, often resulting in incomplete or inaccurate responses. Therefore, tailored for healthcare data, LLaVA-Med [319]was introduced. LLaVA-Med was trained using a novel curriculum learning approach on a large PMC-15M dataset containing 15 million biomedical image-text pairs extracted from PubMed Central. By using GPT-4 to generate instruction-following data from these pairs, the model was trained to handle open-ended conversational tasks about biomedical images. It outperformed previous state-of-the-art models on several benchmarks, such as biomedical visual question answering, highlighting its utility in clinical settings. Other models inspired by LLaVA include [320, 321, 322, 323, 324]
C-F Stable Diffussion
@@
Diffusion models (DMs) [69] can synthesize high-quality images but demand significant computational resources during training. Furthermore, they require costly evaluations in pixel space. Therefore, an efficient model called the stable diffusion model [70] explicitly separates the compressive and generative learning phases. The proposed model employs an autoencoder that generates a lower-dimensional representation space and is computationally efficient. It enhances the computational efficiency of sampling, leveraging the inductive bias of UNet-based diffusion models for spatially structured data and providing a multipurpose compression model applicable to various generative tasks. The proposed diffusion model demonstrates competitive performance in tasks such as image synthesis, inpainting, and super-resolution and reduces computational costs compared to pixel-based DMs. An improved version of the stable diffusion model, known as SDXL [71], incorporates a larger UNet backbone (2.6B parameters in the UNet), additional attention blocks, and a broader cross-attention context.
C-G Industrial pretrained models
@@
Industrial-scale pre-trained models are extensively employed across a spectrum of industrial applications. Here we briefly mention some SOTA models widely adopted in various industries including healthcare. Language Models for Dialog Applications (LaMDA) [325] is built on a transformer architecture that contains 137B parameters tailored for dialog applications. It addresses two challenges faced by language models: safety and factual grounding. The safety enhancements involve filtering responses using a classifier fine-tuned with crowd worker-annotated data, aligning model behavior with human values based on Google AI principles [^7] https://ai.google/responsibility/principles/. Factual grounding enables the model to reference external knowledge sources to ensure responses are grounded in verifiable information.
@@
A more robust model than GPT-3 known as GLaM (Generalist Language Model) [326] has been proposed by Google, which has 1.2 trillion parameters. GLaM achieved better performance on 29 NLP tasks in nearly all zero-shot, one-shot, and few-shot learning scenarios, consuming only one-third of the energy of GPT-3. A more refined version, the Pathways Language Model (PaLM) [301], with 540 billion parameters, has also been presented by Google. It is a decoder-only-style transformer model with multiple modifications, such as SwiGLU activations [315] for MLP intermediate layers, parallel formulation for a faster training speed, multi-query attention, and RoPE embeddings [316] for improved performance on long sequences. PaLM has achieved SOTA performance in reasoning tasks and shows proficiency in multilingual tasks and source code generation. A more efficient model, PaLM 2 [327], was recently developed by Google and performs better than PaLM on various downstream tasks, with enhanced multilingual, reasoning, and coding capabilities. It is utilized at various sizes in 25 Google products, including Bard  [^8] https://bard.google.com/, the health-focused Med-PaLM [50], Med-Plam 2 [226], and the security-focused Sec-PaLM [^9] https://cloud.google.com/security/ai/. SenseNova [^10] https://www.sensetime.com/en/news-detail/51166397?categoryId=1072, developed by SenseTime, is a large-scale FM that offers various API interfaces and services specifically designed for enterprise customers and widely adopted in different industries, mainly in China, such as in the automotive sector, where SenseNova was instrumental in the creation of the BEV (Bird’s-Eye-View) system that can recognize 3,000 types of objects. Furthermore, SenseNova introduced the “SenseChat” [^11] https://chat.sensetime.com/wb/login model, which is similar to ChatGPT but specifically designed for use in China.
TABLE IV: Flagship Foundation Models. E-WP: English Wikipedia, CC: Common Crawl, WT: WebText. The link is provided if the models publicly available.
@@
[13] 2018 BERT 340M [311] , E-WP NLP
[14] 2019 RoBERTa 355M [311], E-WP, CC-News, Open WT [328], stories [329] NLP
[309] 2019 DistilBERT 66M [311], E-WP NLP
[308] 2020 ALBERT 235M E-WP,[311] NLP
@@
[57] 2019 GPT-2 1.5B WT NLP
[28] 2020 GPT-3 175B CC (filtered), WT2, Books1, Books2, E-WP NLP
@@
[312] 2022 InstructGPT 1.3B CC (filtered), WT2, Books1, Books2, E-WP NLP
@@
[314] 2023 LLaMA 65B CC, C4 [16], Github, E-WP, Books, ArXiv, StackExchange NLP
@@
[71] 2023 SDXL 3.5B - Image synthesis
@@
[326] 2022 GLaM 137B Filtered Webpages, E-WP, Conversations, Forums, Books, News NLP tasks